Find the slope , intercept and what to plot on the following equation which also need to be manipulated.

1.y=ax+1/2bx^2

2. Y^2 = b^2 +2ax

3. a/x+b/y = c


1
Expert's answer
2021-03-26T11:38:05-0400

1) Intercept:

"ax+0.5bx^2=0\\x=-\frac{2a}{b}"

Slope:


"k(x)=a+bx"

2) Intercept:

"b^2 +2ax=0\\x=-\frac{b^2}{2a}"

Slope


"k=2a"

3) Slope: none


"k=-\frac{a}{b}"